# Accelerating Quantum Monte Carlo Calculations with Set-Equivariant Architectures and Transfer Learning

### v2

@@ 2.4 Data Generation and Training Procedures
Figure 3: Accuracy of the predictions with and without transfer learning for $N=100$ and $\alpha=2.5$. The standard error over the validation set is computed as $\delta m^{2}_{z}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}\left(m_{z,i}^{2}-\tilde{m}_{z,i}^{2}\right)^{2}}$ where $n$ is the number of data points in that set, $m_{z,i}^{2}$ are the fluctuations of the magnetization calculated via QMC with original variational states and $2048$ samples, and $\tilde{m}_{z,i}^{2}$ are the set-transformer predictions for the same quantities.
The second stage in applying the regressor model consists in analyzing the behavior of the expected values of both the second moment of the magnetization and the entanglement entropy in the vicinity of the critical point. To this end, we generate new reference ground states near the phase transition for each system with $\alpha\in\{1.66,1.7,2.5,6\}$, considering spin chains of lengths $50$, $100$, and $150$ [30]. This finer grid allows us to capture the intricate features of the ground state and the order parameter across the transition, which constitute the most challenging regions of the parameter space. For training, we only use the states corresponding to the extreme values $\alpha\in\{1.66,6\}$ and apply data augmentation to enrich the dataset. For both observables, optimization was performed using stochastic gradient descent with a polynomial learning rate schedule.
@@ 2.5 Transfer Learning
Figure 4: Results of the nested binary classification of states into paramagnetic, ferromagnetic and antiferromagnetic categories for a system of $50$ spins throughout the ranges of $\alpha$ and $J$. This model was trained on $\alpha\in\left\{0,3.2,6\right\}$, highlighted with black rectangles, with up to $1024$ samples per training point. Also shown are the phase boundaries extracted from Refs. [27, 41, 28, 42]
@@ 3 Results
Figure 5: Regression results for the second moment of the magnetization for a system of $50$ spins throughout the ranges of $\alpha$ and $J$. Left: $\langle m^{2}\rangle$ predicted by the set transformer. Right: difference between the predictions and the results obtained via QMC calculations, in logarithmic scale. This model was trained on $\alpha\in\left\{0,3.2,6\right\}$, highlighted with black rectangles, with up to $1024$ samples per training point. Also shown are the phase boundaries extracted from Refs. [27, 41, 28, 42].
We first present the results of the phase-detection classifier in Figure 4, which also depicts the phase boundaries proposed in Refs. [27, 41, 28, 42], color coded according to their source. Despite being trained on only three values of $\alpha$, the classifier works well across the whole spectrum of $\alpha$. The border between the ferro- and paramagnetic regions is particularly well defined, while a few artifacts are visible in the case of the threshold between para- and antiferromagnetic ground-state configurations. Those may be partially attributable to a problem in the training data. The erroneous predictions are well separated from the threshold itself and can easily be identified as such.
@@
Figure 6: Regression results for the fourth moment of the magnetization for a system of $50$ spins throughout the ranges of $\alpha$ and $J$. Left $\langle m^{4}\rangle$ predicted by the set transformer. Right difference in logarithmic scale between the predictions and the results obtained via QMC calculations. This model is trained on $\alpha\in\left\{0,3.2,6\right\}$, highlighted with black rectangles, with up to $1024$ samples per training point. Also shown are the phase boundaries extracted from Refs. [27, 41, 28, 42].
As explained in the previous section, for the second Rényi entanglement entropy, the set transformer was trained on a set of states where each pair $\{\alpha,J/h\}$ has five realizations, but each one of them is labeled by their mean entanglement entropy, instead of the individual values. This is done with the intention of averaging out the numerical fluctuations that arise from the QMC calculations, and at the same time enables us to get averaged outputs of a single state used as input once the regressor is trained. Furthermore, we employ a transfer-learning technique, as explained in Section 2.5, in order to reduce the training cost for the larger system sizes once the model has been trained on the shortest chain. The left panel of Figure 7 shows the predictions of the set transformer for the states of $\alpha=2.5$ for system sizes of $N=50,100,150$, where we also used a Savitzky-Golay filter [43], which allows us to better capture the trend with increased size as we approach the critical point. To further understand the increase of the entanglement with the system size at the phase transition, we study their relationship. In the right panel of Figure 7 we show the predicted logarithmic volume law calculated with the results of the set transformer. This result can be compared with those found in the literature, such as in [30], where the volume law reported was $\langle S_{2}\rangle=0.19\log(N)+0.01$.
Figure 7: Left: dependence of the Rényi 22-entropy on the $J$ parameter for three different chain sizes and a common value of $\alpha=2.5$. The vertical dotted line represents the critical point of the system, which marks the transition from a ferromagnetic ordering to a paramagnetic one. This model is trained on $\alpha\in\left\{1.66,6\right\}$ with up to $1024$ samples per training point. Note that the results presented in this graph correspond to the predictions for a single data point, as opposed to the mean calculation over multiple realizations of each $\{\alpha,J/h\}$ pair. Right: variation of $\langle S_{2}\rangle$ with the system size at the phase transition for $\alpha=2.5$. The black circles show the set-transformer predictions, the black line is the linear regression against these points and the red line is the results from [30]
@@
where $\nu$ and $\beta$ are the standard critical exponents for the divergence of the correlation length and the order parameter, respectively. The function $f$ is a universal scaling function accounting for finite-size effects. To extract the values of $J_{c}$, $\nu$, and $\beta$, we perform a data collapse analysis using the Python library fssa [44, 45].

### v3

@@ 2.4 Data Generation and Training Procedures
Figure 3: Accuracy of the predictions with and without transfer learning for $N=100$ and $\alpha=2.5$. The standard error over the validation set is computed as $\delta m^{2}_{z}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}\left(m_{z,i}^{2}-\tilde{m}_{z,i}^{2}\right)^{2}}$, where $n$ is the number of data points in that set, $m_{z,i}^{2}$ are the fluctuations of the magnetization calculated via QMC with original variational states and $2048$ samples, and $\tilde{m}_{z,i}^{2}$ are the set-transformer predictions for the same quantities.
The second stage in the application of the regressor model consists in analyzing the behavior of the expected values of both the second moment of the magnetization and the entanglement entropy in the vicinity of the critical point. To this end, we generate new reference ground states near the phase transition for each system with $\alpha\in\{1.66,1.7,2.5,6\}$, considering spin chains of lengths $50$, $100$, and $150$ [PhysRevB.110.205147]. This finer grid allows us to capture the intricate features of the ground state and the order parameter across the transition, which constitute the most challenging regions of the parameter space. For training, we only use the states corresponding to the extreme values $\alpha\in\{1.66,6\}$ and apply data augmentation to enrich the dataset. For both observables, optimization was performed using stochastic gradient descent with a polynomial learning rate schedule.
@@ 2.5 Transfer Learning
Figure 4: Results of the nested binary classification of states into paramagnetic, ferromagnetic and antiferromagnetic categories for a system of $50$ spins throughout the ranges of $\alpha$ and $J$. This model was trained on $\alpha\in\left\{0,3.2,6\right\}$, highlighted with black rectangles, with up to $1024$ samples per training point. Also shown are the phase boundaries extracted from Refs. [romanroche2023, zhu2018fidelity, koziol2021quantumcritical, sun2017]
@@ 3 Results
Figure 5: Regression results for the second moment of the magnetization for a system of $50$ spins throughout the ranges of $\alpha$ and $J$. Left: $\langle m^{2}\rangle$ predicted by the set transformer. Right: difference between the predictions and the results obtained via QMC calculations, in logarithmic scale. This model was trained on $\alpha\in\left\{0,3.2,6\right\}$, highlighted with black rectangles, with up to $1024$ samples per training point. Also shown are the phase boundaries extracted from Refs. [romanroche2023, zhu2018fidelity, koziol2021quantumcritical, sun2017].
We first present the results of the phase-detection classifier in Figure 4, which also depicts the phase boundaries proposed in Refs. [romanroche2023, zhu2018fidelity, koziol2021quantumcritical, sun2017], color coded according to their source. Despite being trained on only three values of $\alpha$, the classifier works well across the whole spectrum of $\alpha$. The border between the ferro- and paramagnetic regions is particularly well defined, while a few artifacts are visible in the case of the threshold between para- and antiferromagnetic ground-state configurations. Those may be partially attributable to a problem in the training data. The erroneous predictions are well separated from the threshold itself and can easily be identified as such.
@@
Figure 6: Regression results for the fourth moment of the magnetization for a system of $50$ spins throughout the ranges of $\alpha$ and $J$. Left $\langle m^{4}\rangle$ predicted by the set transformer. Right difference in logarithmic scale between the predictions and the results obtained via QMC calculations. This model is trained on $\alpha\in\left\{0,3.2,6\right\}$, highlighted with black rectangles, with up to $1024$ samples per training point. Also shown are the phase boundaries extracted from Refs. [romanroche2023, zhu2018fidelity, koziol2021quantumcritical, sun2017].
As explained in the previous section, for the second Rényi entanglement entropy, the set transformer was trained on a set of states where each pair $\{\alpha,J/h\}$ has five realizations, but each one of them is labeled by their mean entanglement entropy, instead of the individual values. This is done with the intention of averaging out the numerical fluctuations that arise from the QMC calculations, and at the same time enables us to get averaged outputs of a single state used as input once the regressor is trained. Furthermore, we employ a transfer-learning technique, as explained in Section 2.5, in order to reduce the training cost for the larger system sizes once the model has been trained on the shortest chain. The left panel of Figure 7 shows the predictions of the set transformer for the states of $\alpha=2.5$ for system sizes of $N=50,100,150$, where we also used a Savitzky-Golay filter [Savitzky1964], which allows us to better capture the trend with increased size as we approach the critical point. To further understand the increase of the entanglement with the system size at the phase transition, we study their relationship. In the right panel of Figure 7 we show the predicted logarithmic volume law calculated with the results of the set transformer. This result can be compared with those found in the literature, such as in [PhysRevB.110.205147], where the volume law reported was $\langle S_{2}\rangle=0.19\log(N)+0.01$.
Figure 7: Left: dependence of the Rényi 22-entropy on the $J$ parameter for three different chain sizes and a common value of $\alpha=2.5$. The vertical dotted line represents the critical point of the system, which marks the transition from a ferromagnetic ordering to a paramagnetic one. This model is trained on $\alpha\in\left\{1.66,6\right\}$ with up to $1024$ samples per training point. Note that the results presented in this graph correspond to the predictions for a single data point, as opposed to the mean calculation over multiple realizations of each $\{\alpha,J/h\}$ pair. Right: variation of $\langle S_{2}\rangle$ with the system size at the phase transition for $\alpha=2.5$. The black circles show the set-transformer predictions, the black line is the linear regression against these points and the red line is the results from [PhysRevB.110.205147]
@@
where $\nu$ and $\beta$ are the standard critical exponents for the divergence of the correlation length and the order parameter, respectively. The function $f$ is a universal scaling function accounting for finite-size effects. To extract the values of $J_{c}$, $\nu$, and $\beta$, we perform a data collapse analysis using the Python library fssa [fssa, melchert2009autoscalepy].